# Comparison of Probabilistic Structural Reliability Methods for Ultimate Limit State Assessment of Wind Turbines

### v1

@@ Abstract
The probabilistic design of offshore wind turbines aims to ensure structural safety in a cost-effective way. This involves conducting structural reliability assessments for different design options and considering different structural responses. There are several structural reliability methods, and this paper will apply and compare different approaches in some simplified case studies. In particular, the well known environmental contour method will be compared to a more novel approach based on sequential sampling and Gaussian processes regression for an ultimate limit state case study. For one of the case studies, results will also be compared to results from a brute force simulation approach. Interestingly, the comparison is very different from the two case studies. In one of the cases the environmental contours method agrees well with the sequential sampling method but in the other, results vary considerably. Probably, this can be explained by the violation of some of the assumptions associated with the environmental contour approach, i.e. that the short-term variability of the response is large compared to the long-term variability of the environmental conditions. Results from this simple comparison study suggests that the sequential sampling method can be a robust and computationally effective approach for structural reliability assessment.
@@ 1 Introduction
In this study, the environmental conditions are described by two joint probability distributions of wind speed and turbulence, and a simplified 2-dimensional case study is considered. For one of the case studies, a joint distribution is fitted to wind data for a location near South Brittany, as outlined in Vanem et al. (2023). In additional, an alternative joint model is assumed for a somewhat different wind climate, without reference to any particular location, but henceforth simply referred to as Site A. In both cases, extreme responses of maximum flapwise blade root bending moment are considered for a floating (South Brittany) and a floating (Site A) offshore wind turbine, respectively. In both case studies, a simplified problem is addressed in order to compare the methods, assuming only two environmental input variables; mean wind speed and turbulence at hub height.
@@
In order to alleviate this, two fundamental approaches can be taken: 1) more efficient short-term response analysis and 2) the need for fewer short-term response calculations. In this study, two approaches for long-term extreme response estimation will be compared, i.e. environmental contours and sequential sampling, which both aim at solving the ULS problem with fewer short term response calculations. However, with both approaches a surrogate model is established for making the short-term response analysis more efficient.
@@
Under the assumption that the largest response occurs in the most severe environmental conditions, one may define a limit state function as a function of long-term variables only and construct environmental contours in the long-term variable space. One may then assume that any design with a limit state function fully outside of the environmental contour has a failure probability smaller than the exceedance probability associated with that contour. Alternatively, one may assume that the long-term extreme response with a return period corresponding to the exceedance probability of the environmental contour is the maximum response evaluated at points along the environmental contours. Hence, environmental contours can be used to estimate extreme long-term responses for desired return periods, with only a limited set of short-term response analyses. This approach assumes that extreme long-term response is dominated by the long-term variability of environmental condition, and ignores the short-term variability of the response, but this may be accounted for by using a higher quantile of the conditional short-term extreme response.
@@
The environmental contour method for structural reliability is a well-adopted method, in particular in ocean engineering applications, and it is recommended in DNV’s recommended practice on environmental conditions and environmental loads DNV (2021). Both IFORM (inverse first-order reliability method) Winterstein et al. (1993); Haver and Winterstein (2009) and DS (direct sampling) Huseby et al. (2013, 2015) contours are applied to calculate 50-year contours at site A. The output quantity of interest (QoI) considered here is the maximum flapwise blade root bending moment, and the contour method is used to estimate the long-term extreme maximum flapwise blade root bending moment. The different contour methods yield slightly different results, but are generally comparable. Similar results are obtained using a Gaussian process regression and sequential sampling approach Gramstad et al. (2020), and also compared with brute force simulation results. The results from the contour methods are generally in line with brute force results if the 90% fractile of the short-term extreme response distribution is used.
@@ 1.1 Problem statement and definitions
The goal of this paper is to compare different methods for ULS reliability assessment of offshore wind structures. A simplified case with two environmental input variables is defined, where the response of interest is the maximum flapwise blade root bending moment. The environmental contour method and a sequential sampling method are applied, that may serve as a benchmark to be compared with other approaches in future case studies.
@@
The environmental contours approach is a well-established method, that is recommended in DNV’s recommended practice DNV (2021). Different variants of this method exist Haver and Winterstein (2009); Huseby et al. (2013); Leira (2008); Silva-González et al. (2013); Montes-Iturrizaga and Heredia-Zavoni (2015); Li et al. (2016); Lutes and Winterstein (2016); Haselsteiner et al. (2017); Chai and Leira (2018); Manuel et al. (2018); Ross et al. (2019); Haselsteiner et al. (2021), and several applications of this approach are reported in the literature, especially ocean engineering applications Niedzwwecki et al. (1998); Baarholm and Moan (2001); Baarholm et al. (2010); Muliawan et al. (2013); Vanem (2018), but also for offshore wind Velarde et al. (2019), earthquake engineering van de Lindt and Niedzwecki (2000) and bridge design Giske et al. (2018); Castellon et al. (2022). The Gaussian processes surrogate model with sequential sampling approach is a more recent development Mohamad and Sapsis (2018); Gramstad et al. (2020), with perhaps less experience although several applications have been reported with such approaches to structural reliability problems as well Echard et al. (2011); Agrell and Dahl (2021); Gond and Pan (2022); Castellon et al. (2023); Lv et al. (2015); Sun et al. (2017); Lelièvre et al. (2018); Peng et al. (2024); see also the recent overviews in Teixeira et al. (2021); Moustapha et al. (2022). This paper outlines a study where these methods are applied to two offshore wind case studies of ULS reliability assessment, as described in Cousin et al. (2023).
@@
Response calculations for different environmental conditions are performed with Hipersim (https://gitlab.windenergy.dtu.dk/HiperSim/hipersim), used to generate turbulence wind boxes, and a fast-to-evaluate surrogate for the NREL 5MW reference wind turbine Jonkman et al. (2009) developed by ETH Zürich Dimitrov et al. (2022); Schär et al. (2023), used to calculate the wind turbine responses.
@@
Two case studies are considered, with offshore wind turbines at two different locations, i.e. Site A and South of Brittany. For both cases, joint statistical models for the relevant input variables were taken as given, but the model for the South Brittany site has been fitted to wind and wave data as outlined in Vanem et al. (2023). Note, however, that the statistical models were initially fitted to higher-dimensional data, but only mean wind speed and turbulence are considered in this study.
@@
This distribution is fitted to hindcast data from the ANEMOC (Digital Atlas of Ocean and Coastal Sea States) database [^1] URL: http://anemoc.cetmef.developpement-durable.gouv.fr/, covering 32 years of data from the years 1979 to 2010. It should be noted that these data do not include the turbulence variable, $\sigma_{U}$, and therefore a conditional distribution for turbulence, conditioned on mean wind speed, was established based on the normal turbulence model described in IEC (2005), i.e., a conditional lognormal distribution as also outlined in Vanem et al. (2023).
It should be noted that the ETH surrogate model has been trained on the NREL 5 MW turbine whose hub is at 90 m, and the mean wind speed U is not translated to the hub height 90 m in this exercise. Hence, the actual response estimates cannot be used directly, but this is deemed appropriate for a comparison exercise, as long as the same wind input is used in all cases that are to be compared. The response model is only applicable between the cut-in speed (3 m/s) and cut-out speed (25 m/s), so sets of random variables with mean wind speed $U$ below cut-in wind speed (3 m/s) or above cut-out wind speed (25 m/s) were discarded. However, these conditions above the cut-out wind speed are not assumed to contribute to the long-term extreme blade root responses for these turbines, so this is deemed reasonable.
@@ 2.1 Computational model
The computational model considered in this study is the well-known NREL 5MW onshore wind turbine Jonkman et al. (2009) equipped with NREL’s ROSCO (Reference Open-Source Controller) Abbas et al. (2022) and simulated with the open-source aero-servo-elastic simulator code OpenFAST NREL (2021).
@@ 3.1 Site A
The points on the 50-year return period contours are taken as input for the surrogate model. There are 68 input points from the 50-year DS-contour, and 40 input points from the 50-year IFORM contour. In total 1000 seeds are run with mNARX (10-minute simulation) for each input point on both the DS-contour and IFORM contour. However, points corresponding to wind conditions below the cut-in wind speed (3 m/s) and above the cut-out wind speed (25 m/s) are disregarded. The response model is not applicable to these conditions, and it is tacitly assumed that these conditions will not contribute to the long-term extreme response. Hence, for wind conditions outside the operational range of the wind turbine, the response is simply set to zero. It is noted that the mNARX model sometimes gives NaN (not a number) for certain seeds for some wind conditions. This is most likely due to negative or extremely high instantaneous winds speeds since this model is not trained neither at negative wind speed nor at extremely high wind speeds. This could occur for example for a combination of high mean wind speed and high turbulence. In such cases, another seed is simply selected to avoid that mNARX gives NaN.
@@ 4.2.2 Parametric model for the short-term response
In the present work two different models $\hat{g}_{Y|\bm{X}}(y|\boldsymbol{\theta}(\bm{x}))$ for the short-term extreme response have been considered: the Gumbel-distribution and the generalized extreme value (GEV) distribution. The Gumbel distribution has two parameters $\boldsymbol{\theta}=(\alpha,\beta)$ (location and scale) and the GEV distribution has three parameters $\boldsymbol{\theta}=(\alpha,\beta,\gamma)$ (location, scale and shape).
@@ 4.2.3 GP-model for distribution parameters
Given some training data, i.e. observed distribution parameters $\boldsymbol{\zeta}_{j}=(\zeta_{j}^{(1)},\cdots,\zeta_{j}^{(m)})$ for points $\bm{x}_{j}$: $D=\{(\bm{x}_{j},\boldsymbol{\zeta}_{j}\}_{j=1}^{N}$ one can derive the posterior predictive distribution for "new" points under the observed training data. As described in section 4.2.2 the distribution parameters are assumed to come with Gaussian noise, so that $\boldsymbol{\zeta}_{j}=\theta(\bm{x}_{j})+N(0,\Sigma_{j})$, where $\Sigma_{j}$ is the covariance matrix of each set of the $m$ distribution parameters, as estimated using the procedure described in section 4.2.2. For further details on the actual implementation it is referred to the Python source code Agrell and Gramstad (2023).

### v2

@@ Abstract
The probabilistic design of offshore wind turbines aims to ensure structural safety in a cost-effective way. This involves conducting structural reliability assessments for different design options and considering different structural responses. There are several structural reliability methods, and this paper will apply and compare different approaches in some simplified case studies. In particular, the well known environmental contour method will be compared to a more novel approach based on sequential sampling and Gaussian processes regression for an ultimate limit state case study on the maximum flapwise blade root bending moment. For one of the case studies, results will also be compared to results from a brute force simulation approach. Interestingly, the comparison is very different from the two case studies. In one of the cases the environmental contours method agrees well with the sequential sampling method but in the other, results vary considerably. Probably, this can be explained by the violation of some of the assumptions associated with the environmental contour approach, i.e. that the short-term variability of the response is large compared to the long-term variability of the environmental conditions. Results from this simple comparison study suggests that the sequential sampling method can be a robust and computationally effective approach for structural reliability assessment.
@@ 1 Introduction
In this study, the environmental conditions are described by two joint probability distributions of wind speed and turbulence, and a simplified 2-dimensional case study is considered. For one of the case studies, a joint distribution is fitted to wind data for a location near South Brittany, as outlined in Vanem et al. (2023, 2024). In addition, an alternative joint model is assumed for a somewhat different wind climate, without reference to any particular location, but henceforth simply referred to as Site A. In both cases, extreme responses of maximum flapwise blade root bending moment are considered for a floating (South Brittany) and a bottom-fixed (Site A) offshore wind turbine, respectively. In both case studies, a simplified problem is addressed in order to compare the methods, assuming only two environmental input variables; mean wind speed and turbulence at hub height. Hence, the contributions to the dynamic response from wave excitation are not included in the models, which essentially uses a generic onshore wind turbine placed in two offshore environments.
@@
In order to alleviate this, two fundamental approaches can be taken: 1) more efficient short-term response analysis and 2) the need for fewer short-term response calculations. In this study, two approaches for long-term extreme response estimation will be compared, i.e. environmental contours and sequential sampling, which both aim at solving the ULS problem with fewer short term response calculations. However, with both approaches a surrogate model is established for making the short-term response analysis more efficient, but it is noted that this is not strictly necessary and is not an integral part of any of the methods; it is simply applied in this study as a substitute for more computationally demanding calculation methods to speed up short term response calculations.
@@
Under the assumption that the largest response occurs in the most severe environmental conditions, one may define a limit state function as a function of long-term variables only and construct environmental contours in the long-term variable space. One may then assume that any design with a limit state function fully outside of the environmental contour has a failure probability smaller than the exceedance probability associated with that contour. Alternatively, one may assume that the long-term extreme response with a return period corresponding to the exceedance probability of the environmental contour is the maximum response evaluated at points along the environmental contours. Hence, environmental contours can be used to estimate extreme long-term responses for desired return periods, with only a limited set of short-term response analyses. This approach assumes that extreme long-term response is dominated by the long-term variability of environmental condition, and ignores the short-term variability of the response (the conditional variability of the response in an environmental condition), but this may be accounted for by using a higher quantile of the conditional short-term extreme response. A comparison between environmental contours and response-based approaches to reliability assessment of floating structures is presented in Zhao and Dong (2022).
Reliability-based design methods for extreme response assessment of offshore wind turbines are proposed in Cheng et al. (2003), where environmental contours are used to identify design load cases, and different statistical methods are used to extrapolate conditional extreme responses. The long-term distribution of extreme response is then the convolution of conditional short-term response distributions over the long-term distribution of sea states. Several other studies have focused on how to best combine the variability of the long-term environmental conditions and the short-term conditional extreme responses to extrapolate and estimate long-term extreme responses of wind turbines, see e.g. Moriarty et al. (2004); Fogle et al. (2008); Agarwal and Manuel (a, b); Schinas et al. (2021); Yang et al. (2022). Datasets for for validation of such techniques have been provided in Moriarty . This is very well aligned with the approaches compared in this paper, which aims at comparing two computationally efficient methods for estimating long-term extreme responses by combining joint distributions of long-term conditions with computationally costly conditional response calculations. This is illustrated by simple case studies, where only the wind conditions are considered for the long-term environmental states and hence ignoring the variability of sea states. For bottom-fixed structures this may be fine; the flap moments may not be notably affected by changes in sea states (as also suggested in Cheng et al. (2003)), but it may be an over-simplification for turbines with floating support structures where wave- and wind-induced surge and pitch motions could give rise to inertia forces on the blades (see e.g. Tran et al. (2015); Tran and Kim (2015); Lamei and Hayatdavoodi (2020)). Notwithstanding, the methods for ULS assessment presented in this paper could be extended to also account for long-term variations in sea states, and the simplified case study is still believed to provide a useful relative comparison between the environmental contour and the sequential sampling methods.
It should be noted that the fundamental assumption underpinning the environmental contour method may not be fulfilled for many offshore wind turbine responses, where the structural responses might be non-monotonic due to operational constraints or active and passive structural control Abdallah et al. (2016). In order to account for this, a modified environmental contour method was proposed in Li et al. (2016), which considers additional contours within different operational regimes, i.e. operational or parked. This approach has later been applied to various offshore wind problems, see e.g. Li et al. (2017, 2019); Lee et al. (2023), and an extended method was presented in Chen et al. (2020). A comparison study presented in Li et al. (2018) indicates that the conventional environmental contour method might under-estimate extreme wind-dominated responses compared to the modified contour method. However, in that study, turbulence was not included as an environmental variable, and it was shown in Chen et al. (2020) that including turbulence as a stochastic variable improved results with the original contour method significantly. Indeed, it was found that by including turbulence as an additional stochastic environmental variable the original environmental contour method is able to predict extreme responses from wind-induced loads satisfactorily. In the study presented in this paper, turbulence is indeed included as a stochastic variable, and only the original environmental contour method is included in the comparison study.
@@
The environmental contour method for structural reliability is a well-adopted method, in particular in ocean engineering applications, and it is recommended in DNV’s recommended practice on environmental conditions and environmental loads DNV (2021) as well as in the NORSOK standard on actions and action effects NORSOK (2017). Both IFORM (inverse first-order reliability method) Winterstein et al. (1993); Haver and Winterstein (2009) and DS (direct sampling) Huseby et al. (2013, 2015) contours are applied to calculate 50-year contours at site A. The output quantity of interest (QoI) considered here is the maximum flapwise blade root bending moment, and the contour method is used to estimate the long-term extreme maximum flapwise blade root bending moment. The different contour methods yield slightly different results, but are generally comparable. Similar results are obtained using a Gaussian process regression and sequential sampling approach Gramstad et al. (2020), and also compared with brute force simulation results. The results from the contour methods are generally in line with brute force results if the 90% fractile of the short-term extreme response distribution is used.
@@ 1.1 Problem statement and definitions
The goal of this paper is to compare different methods for ULS reliability assessment of offshore wind structures. A simplified case with two environmental input variables is defined, where the response of interest is the maximum flapwise blade root bending moment. The environmental contour method and a sequential sampling method are applied, that may serve as a benchmark to be compared with other approaches in future case studies. It is acknowledged that the 2D case that only considers the variability of the long-term wind conditions is a simplification and the effect of wave induced motions are ignored. Although this might be reasonable for the bottom-fixed structure, it is more questionable for the floating case. Nevertheless, it is believed that the comparison study between the different approaches is still useful, since the same assumptions are made for both approaches.
@@
The environmental contours approach is a well-established method, that is recommended in DNV’s recommended practice DNV (2021). Different variants of this method exist Haver and Winterstein (2009); Huseby et al. (2013); Leira (2008); Silva-González et al. (2013); Montes-Iturrizaga and Heredia-Zavoni (2015); Li et al. (2016); Lutes and Winterstein (2016); Haselsteiner et al. (2017); Chai and Leira (2018); Manuel et al. (2018); Ross et al. (2019); Haselsteiner et al. (2021); de Hauteclocque et al. (2022), and several applications of this approach are reported in the literature, especially ocean engineering applications Niedzwwecki et al. (1998); Baarholm and Moan (2001); Baarholm et al. (2010); Muliawan et al. (2013); Vanem (2018); Liao et al. (2022), but also for offshore wind turbines Saranyasoontorn and Manuel (2004, 2005, 2006); Valamanesh et al. (2015); Velarde et al. (2019); Raed et al. (2020); Katsikogiannis et al. (2024), earthquake engineering van de Lindt and Niedzwecki (2000), bridge design Giske et al. (2018); Castellon et al. (2022) and ice loads Chai et al. (2020). The Gaussian processes surrogate model with sequential sampling approach is a more recent development Mohamad and Sapsis (2018); Gramstad et al. (2020), with perhaps less experience although several applications have been reported with such approaches to structural reliability problems as well Echard et al. (2011); Agrell and Dahl (2021); Gond and Pan (2022); Castellon et al. (2023); Lv et al. (2015); Sun et al. (2017); Lelièvre et al. (2018); Peng et al. (2024); see also the recent overviews in Teixeira et al. (2021); Moustapha et al. (2022). This paper outlines a study where these methods are applied to two offshore wind case studies of ULS reliability assessment, as described in Cousin et al. (2023).
@@
Response calculations for different environmental conditions are performed with Hipersim (https://gitlab.windenergy.dtu.dk/HiperSim/hipersim), used to generate turbulence wind boxes assuming the Mann turbulence spectrum, and a fast-to-evaluate surrogate for the NREL 5MW reference wind turbine Jonkman et al. (2009) developed by ETH Zürich Dimitrov et al. (2022); Schär et al. (2024), used to calculate the wind turbine responses. It is noted that this surrogate (to be further detailed in section 2 below) is only used as a faster wind turbine simulator and that it is not an integral part of the approaches compared in this study. It is useful to speed up response calculations in this methodological study, but it is noted that the approaches to long-term extreme response assessment would not be dependent on this surrogate.
@@
Two case studies are considered, with offshore wind turbines at two different locations, i.e. Site A and South of Brittany. For both cases, joint statistical models for the relevant input variables were taken as given, but the model for the South Brittany site has been fitted to wind and wave data as outlined in Vanem et al. (2023, 2024). Note, however, that the statistical models were initially fitted to higher-dimensional data, but only mean wind speed and turbulence are considered in this study.
@@
This distribution is fitted to hindcast data from the ANEMOC (Digital Atlas of Ocean and Coastal Sea States) database [^1] URL: http://anemoc.cetmef.developpement-durable.gouv.fr/, covering 32 years of data from the years 1979 to 2010. It should be noted that these data do not include the turbulence variable, $\sigma_{U}$, and therefore a conditional distribution for turbulence, conditioned on mean wind speed, was established based on the normal turbulence model described in IEC (2005), i.e., a conditional lognormal distribution as also outlined in Vanem et al. (2023). For the mean wind speed, a marginal hybrid distribution consisting of a Weibull distribution for the body and a generalized Pareto distribution for the tails was assumed.
It should be noted that the ETH surrogate model has been trained on the NREL 5 MW turbine whose hub is at 90 m, and the mean wind speed U is not translated to the hub height 90 m in this exercise. Hence, the actual response estimates cannot be used directly, but this is deemed appropriate for a comparison exercise, as long as the same wind input is used in all cases that are to be compared. The response model is only applicable between the cut-in speed (3 m/s) and cut-out speed (25 m/s), so sets of random variables with mean wind speed $U$ below cut-in wind speed (3 m/s) or above cut-out wind speed (25 m/s) were discarded. However, these conditions above the cut-out wind speed are not assumed to contribute to the long-term extreme blade root responses for these turbines, since the turbine will then be in parked conditions, so this is deemed reasonable.
@@ 2.1 Computational model
The computational model considered in this study is the well-known NREL 5MW onshore wind turbine Jonkman et al. (2009) equipped with NREL’s ROSCO (Reference Open-Source Controller) Abbas et al. (2022) and simulated with the open-source aero-servo-elastic simulator code OpenFAST NREL (2021). The cut-in, rated and cut-out wind-speeds for this turbine are 3 m/s, 11.4 m/s and 25 m/s, respectively. Further details about this turbine, including natural frequencies and graphs for steady-state responses as a function of wind speed, can be found in Jonkman et al. (2009).
@@ 2.2.2 Manifold construction
Regarding the physically meaningful auxiliary quantities, note that mNARX is a multi-stage surrogate model. It does not directly predict the quantity of interest (QoI) (the blade moment), but first constructs variables that are related to the dynamics of this quantity. These variables can be extracted from the raw input (the wind) or constructed by using another NARX model. They are usually easier to emulate than the main QoI and serve as additional inputs for main NARX model that predicts the QoI. Examples of auxiliary quantities constructed in this study are the blade pitch control system, rotor orientation and rotor speed. For more details of the mNARX surrogate, reference is made to Schär et al. (2024).
@@ 3.1 Site A
The points on the 50-year return period contours are taken as input for the surrogate model. There are 68 input points from the 50-year DS-contour, and 40 input points from the 50-year IFORM contour. In total 1000 seeds are run with mNARX (10-minute simulation) for each input point on both the DS-contour and IFORM contour. However, points corresponding to wind conditions below the cut-in wind speed (3 m/s) and above the cut-out wind speed (25 m/s) are disregarded. The response model is not applicable to these conditions, and it is tacitly assumed that these conditions will not contribute to the long-term extreme response. Hence, for wind conditions outside the operational range of the wind turbine, the response is simply set to zero. It is noted that the mNARX model sometimes gives NaN (not a number) for certain seeds for some wind conditions. This is most likely due to negative or extremely high instantaneous winds speeds since this model is not trained neither at negative wind speed nor at extremely high wind speeds. This could occur for example for a combination of high mean wind speed and high turbulence. In such cases, another seed is simply selected to avoid that mNARX gives NaN. This is not believed to significantly impact the final results if this study, and this additional challenge is outweighed by the fact that utilizing this surrogate makes this comparison study feasible.
@@ 4.2.2 Parametric model for the short-term response
In the present work two different models $\hat{g}_{Y|\bm{X}}(y|\boldsymbol{\theta}(\bm{x}))$ for the short-term extreme response have been considered: the Gumbel-distribution and the generalized extreme value (GEV) distribution. The Gumbel distribution has two parameters $\boldsymbol{\theta}=(\alpha,\beta)$ (location and scale) and the GEV distribution has three parameters $\boldsymbol{\theta}=(\alpha,\beta,\gamma)$ (location, scale and shape). Note that this corresponds to the MAX approach according to Cheng et al. (2003), where only maxima are used to extrapolate the conditional short-term response distributions; see also Ragan and Manuel (2008); Toft et al. (2011); Vanem (2015) for a discussion on different methods for extrapolating short-term conditional extreme responses.
@@ 4.2.3 GP-model for distribution parameters
Given some training data, i.e. observed distribution parameters $\boldsymbol{\zeta}_{j}=(\zeta_{j}^{(1)},\cdots,\zeta_{j}^{(m)})$ for points $\bm{x}_{j}$: $D=\{(\bm{x}_{j},\boldsymbol{\zeta}_{j})\}_{j=1}^{N}$ one can derive the posterior predictive distribution for "new" points under the observed training data. As described in section 4.2.2 the distribution parameters are assumed to come with Gaussian noise, so that $\boldsymbol{\zeta}_{j}=\theta(\bm{x}_{j})+N(0,\Sigma_{j})$, where $\Sigma_{j}$ is the covariance matrix of each set of the $m$ distribution parameters, as estimated using the procedure described in section 4.2.2. For further details on the actual implementation it is referred to the Python source code Agrell and Gramstad (2023).